# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v1

@@ Abstract
Neural network-based policies have demonstrated success in many robotic applications, but often lack human-explanability, which poses challenges in safety-critical deployments. To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification to describe a robot policy in a interpretable form. Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which do not give meaningful insights into the underlying policy. We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning. We also introduce three novel explainability evaluation metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification metrics. Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy. This work bridges neural network policy learning with formal methods, contributing to safer and more transparent decision-making in robotics.
@@ I Introduction
Recent innovations in learning-based methods, particularly those using neural networks, have significantly improved both high-level decision-making and low-level control in robotics. However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount. To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2, 3], including post-hoc analysis techniques and methods to learn policies that are inherently explainable.
In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy. While existing approaches, like saliency maps [4] and feature attribution [5], provide valuable information, they often lack language-based explanations, which limits human-interpretability. Transformer-based language models have recently shown promising results for providing language-based explanations [6, 7, 8], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [9].
@@
The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data. These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [18], fault diagnosis [19], and weather forecasting [20]. However, existing work only focuses on classification accuracy, which does not fully capture the actual explanability of learned specifications. For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints). Some works address verbosity by incorporating heuristic simplification methods, like only keeping the top-$k$ weighted predicates [18, 20] or using a greedy process to remove weights [21]. However, as we show in our empirical results, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed). Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses. TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained, as shown in Figure 2. This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning. We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduce learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices. We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics. TLNet bridges the gap between neural network-based policy learning and providing formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
@@ II Preliminaries
Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},..,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon. We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
@@ II-A Weighted Signal Temporal Logic
We consider wSTL syntax [17] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a real-valued function $f:S\rightarrow\mathbb{R}$ over $S$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively. For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator. For temporal operators, we define the weights as $w:=[w_{k}]_{k\in I}\in\mathbb{R}^{\lvert I\rvert}_{>0}$ for interval $I$. For notational simplicity, we use the following notation interchangeably: $\bigwedge_{i=1}^{N}{}^{w}\phi_{i}=\bigwedge_{i=1}^{N}w_{i}\phi_{i}$ and $\bigvee_{i=1}^{N}{}^{w}\phi_{i}=\bigvee_{i=1}^{N}w_{i}\phi_{i}$.
@@ II-B Normalized wSTL
Our neural network architecture proposed in Section III-A requires the weights within each operator to be normalized. We discuss the reasons for this requirement in Section III-E. Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=1$ and $\sum_{k\in I}w_{k}=1$, where $0\leq w_{i},w_{k}\leq 1$. We allow the weights to be zero because the zero weights do not influence the overall robustness [18].
@@
$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}).$ (4)
@@
Building from [18], we define the aggregation functions as follows,
@@
$\otimes^{\mathcal{G}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
$\oplus^{\mathcal{F}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (10)
where $\sigma>0$ is a scaling constant.
@@ II-C Problem Statement
We consider the following binary classification task. Let $\mathcal{T}=\Set{(\tau_{i},l_{i})}{i=1,\dots,N_{\mathrm{data}}}$ be a labeled dataset, where $\tau_{i}\in\mathbb{R}^{n\times H}$ is the $i^{th}$ trajectory with label $l_{i}\in\{1,-1\}$ and $N_{\mathrm{data}}\in\mathbb{Z}_{>0}$ is the number of data points in the dataset.
@@ III Methodology
We first introduce our neural network architecture, TLNet, which allows us to represent wSTL structures in a fully differentiable form. We then introduce our simplification process, which enhances the explainability of the generated explanations, and our trajectory sampling process. Finally, we propose novel evaluation metrics for explanability—conciseness, consistency, and strictness—and provide theoretical justification for our network architecture. Algorithm 1 summarizes our overall approach.
@@
1:Dataset $\mathcal{T}$, predicates $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$, similarity threshold $S_{\mathrm{th}}$, number of pruning iterations $N_{\mathrm{pr}}$, number of weights to prune $N_{\mathbf{w}}$, loss function $\tilde{\mathcal{L}}.$
@@
11:end for
@@
III-A Neural Network Architecture
Let $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a predefined set of atomic predicates and $\Psi_{\neg}=\Set{\neg\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a set of their negations. We focus on representing wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
$\phi=0.5\mathcal{F}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]\\ \wedge 0.5\mathcal{G}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
where $w_{ij}^{\mathcal{F}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates defining specification $\phi$. We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [22, 16]. We define the weights $w^{U}$ for temporal operators over interval $I=[0,H]$ to be $w^{U}=[w^{U}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$. We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation.
The architecture of TLNet is designed to ensure the built-in architectural interpretability of the network weights, achieved by the bijective correspondence of the network weights and the importance weights of Equation 11. In Section III-E, we detail the theoretical justification of the formula manipulation to aggregate the weights of the conjunctive and disjunctive clauses in Equation 11. The output of TLNet is the robustness of the input trajectory with respect to the specification defined by the current network weights.
An example TLNet architecture for a dataset containing two predicates is shown in Figure 3. The specification represented by this architecture is $\phi=0.5\mathcal{F}_{I}^{w^{U}}[(w_{11}^{\mathcal{F}}\psi_{1}\vee w_{12}^{\mathcal{F}}\psi_{2}\vee w_{13}^{\mathcal{F}}\neg\psi_{1}\vee w_{14}^{\mathcal{F}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{F}}\psi_{1}\vee w_{22}^{\mathcal{F}}\psi_{2}\vee w_{23}^{\mathcal{F}}\neg\psi_{1}\vee w_{24}^{\mathcal{F}}\neg\psi_{2})]\wedge 0.5\mathcal{G}_{I}^{w^{U}}[(w_{11}^{\mathcal{G}}\psi_{1}\vee w_{12}^{\mathcal{G}}\psi_{2}\vee w_{13}^{\mathcal{G}}\neg\psi_{1}\vee w_{14}^{\mathcal{G}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{G}}\psi_{1}\vee w_{22}^{\mathcal{G}}\psi_{2}\vee w_{23}^{\mathcal{G}}\neg\psi_{1}\vee w_{24}^{\mathcal{G}}\neg\psi_{2})]$.

### v2

@@ Abstract
Learning-based policies have demonstrated success in many robotic applications, but often lack explainability, which poses challenges in safety-critical deployments. To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification which describes a robot policy in a human-interpretable form. Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which limits meaningful insights into the underlying policy. We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning. We also introduce three novel explainability evaluation metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification metrics. Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy. This work bridges policy learning with explainability through formal methods, contributing to safer and more transparent decision-making in robotics.
@@ I Introduction
Recent innovations in learning-based methods, particularly those using neural networks, have significantly improved both high-level decision-making and low-level control in robotics. However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount. To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2], including post-hoc analysis techniques [3, 4] and methods to learn policies that are inherently explainable [5, 6].
In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy. While existing approaches, like saliency maps [7] and feature attribution [8], provide valuable information, they often lack language-based explanations, which limits human-interpretability. Autoregressive language models have recently shown promising results for providing language-based explanations [9, 10, 11], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [12].
@@
The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data. These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [21], fault diagnosis [22], and weather forecasting [23]. However, existing work only focuses on classification accuracy, which does not fully capture the actual explainability of learned specifications. For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints). Some works address verbosity by incorporating heuristic simplification methods, like only keeping the top-$k$ weighted predicates [21, 24] or using a greedy process to remove weights [25]. However, as we show in our empirical results, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed). Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses. TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained. This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning. We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduces learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices. We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics. TLNet bridges the gap between neural network-based policy learning and provides formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
@@ II Preliminaries
Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},\dots,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon of $\tau$. We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
@@ II-A Weighted Signal Temporal Logic
We consider wSTL syntax [20] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a function $f:S\rightarrow\mathbb{R}$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively. For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator. For temporal operators, we define the weights as $w:=[w_{k}]_{k\in I}\in\mathbb{R}^{\lvert I\rvert}_{>0}$ for interval $I$. For notational simplicity, we use the following notation interchangeably: $\bigwedge_{i=1}^{N}{}^{w}\phi_{i}=\bigwedge_{i=1}^{N}w_{i}\phi_{i}$ and $\bigvee_{i=1}^{N}{}^{w}\phi_{i}=\bigvee_{i=1}^{N}w_{i}\phi_{i}$.
@@ II-B Normalized wSTL
Our neural network architecture proposed in Section III-A requires the weights within each operator to be normalized. We discuss the reasons for this requirement in Section III-E. Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=\sum_{k\in I}w_{k}=1$, where $w_{i},w_{k}\in[0,1]$. We allow the weights to be zero because the zero weights do not influence the overall robustness [21].
@@
$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (4)
@@
Building from [21], we define the aggregation functions with a scaling constant $\sigma>0$ as follows,
@@
$\otimes^{\mathcal{G}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
$\oplus^{\mathcal{F}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}.$ (10)
We use $\sigma=0.5$ in this work.
@@ II-C Problem Statement
We consider the following binary classification task. Let $\mathcal{T}=\Set{(\tau_{i},l_{i})}_{i=1:N_{\mathrm{data}}}$ be a labeled dataset, where $\tau_{i}\in\mathbb{R}^{n\times H}$ is the $i^{th}$ trajectory with label $l_{i}\in\{1,-1\}$ and $N_{\mathrm{data}}\in\mathbb{Z}_{>0}$ is the number of data points in the dataset.
@@ III Methodology
We first introduce our neural network architecture, TLNet, which allows us to represent wSTL structures in a fully differentiable form. We then introduce our simplification process, which enhances the explainability of inferred explanations, and our trajectory sampling process. Finally, we propose novel evaluation metrics for explainability—conciseness, consistency, and strictness—and provide theoretical justification for our network architecture. Figure 2 and Algorithm 1 summarize our overall approach.
III-A Neural Network Architecture
Let $\Psi=\Set{\psi_{i}}_{i=1:N_{\mathrm{AP}}}$ be a predefined set of atomic predicates and $\Psi_{\neg}=\Set{\neg\psi_{i}}_{i=1:N_{\mathrm{AP}}}$ be a set of their negations. We focus on representing normalized wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
ϕ=italic-$\phi=$ $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{F}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]$
$\wedge$ $0.5\mathcal{G}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{G}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
where $w_{i}^{\mathcal{F}},w_{ij}^{\mathcal{F}},w_{i}^{\mathcal{G}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates. We define the weights $\bar{w}$ for temporal operators over interval $I=[0,H]$ to be $\bar{w}=[\bar{w}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$. We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation. We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [26, 19].
We represent specifications of the form given in Equation 11 by designing a neural network $f_{\mathbf{w}}$ parameterized by $\mathbf{w}$, where $\mathbf{w}$ defines the importance weights of the currently represented specification $\phi_{\mathbf{w}}$. The input to the neural network is a sequence of the robustness values for each predicate $\psi\in\Psi\cup\Psi_{\neg}$ for each timestep in a trajectory $\tau$. The output is the robustness of the current specification, $r(\phi_{\mathbf{w}},\tau)$, for that trajectory. The neural network is trained to minimize a regularized classification loss (discussed in Section III-B), given a labeled dataset $\mathcal{T}$.
@@ III-A Neural Network Architecture
1:Dataset $\mathcal{T}$, predicates $\Psi=\Set{\psi_{i}}_{i=1:N_{\mathrm{AP}}}$, similarity threshold $S_{\mathrm{th}}$, number of pruning iterations $N_{\mathrm{pr}}$, number of weights to prune $N_{\mathbf{w}}$, loss function $\tilde{\mathcal{L}}.$
@@
11:     if only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$/$\mathbf{w}^{\mathcal{G}}$ then break.
@@
We reduce the number of parameters that are optimized by distributing the weights in Equation 11, resulting in a specification of the following form,
[…]
$0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}}\psi_{j}\right]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}}\psi_{j}\right],$ (12)
where $\mathbf{w}=\mathbf{w}^{\mathcal{F}}\mathbin{\|}\mathbf{w}^{\mathcal{G}}$, $\mathbf{w}^{\mathcal{F}}=[\tilde{w}_{ij}^{\mathcal{F}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\mathbf{w}^{\mathcal{G}}=[\tilde{w}_{ij}^{\mathcal{G}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\tilde{w}_{ij}^{\mathcal{F}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{F}}$, and $\tilde{w}_{ij}^{\mathcal{G}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{G}}$. We justify this weight distribution operation in Section III-E. An example TLNet architecture for a dataset containing two predicates is shown in Figure 3.